OpenStudy (lopezking1):

2(x-4)-12=2-3(x+4) show work

OpenStudy (ivettef365):

First solve what is in parenthesis

OpenStudy (lopezking1):

its not PEMDAS its distributive

OpenStudy (ivettef365):

i know, but you need to do parenthesis first

OpenStudy (luigi0210):

distribute

OpenStudy (anonymous):

use foil 2x -8 -12 =2-3X-12 then solve as u would a normal linear equation

OpenStudy (ivettef365):

right then combine like terms, that is the second step

OpenStudy (lopezking1):

(2x-8)+(-12)=2+-3(x+4)?

OpenStudy (anonymous):

nope 2x+3x=8+12+2-12

OpenStudy (lopezking1):

ohhh i see you reordered it okay

OpenStudy (anonymous):

yess

OpenStudy (lopezking1):

5x=10?

OpenStudy (anonymous):

yep that gives you 2 to check ur answer sub it in the original eqution first on the left...check wat u get then on the right ...check what u get if they are equal ur answer is CORRECT

OpenStudy (lopezking1):

Awesome thanks

OpenStudy (anonymous):

cool dude....!! enjoy equations